OpenStudy (praxer):

In a business meeting, each person shakes hands with each other person, with the exception of Mr. L. Since Mr. L arrives after some people have left, he shakes hands only with those present. If the total number of handshakes is exactly 100, how many people left the meeting before Mr. L arrived? (Nobody shakes hands with the same person more than once.

OpenStudy (rational):

It takes \(2\) persons for one handshake. If \(n\) persons attended the party "excluding" Mr. L and \(k\) of them left before arrival of Mr.L. Then \[\large \binom{n}{2} + n-k = 100\]

OpenStudy (praxer):

I think Mr L was also involved in shaking hands in the above equation it is like you include only the initially present people.

OpenStudy (rational):

Mr L was not counted in first "n" persons' handshakes

OpenStudy (rational):

By the time Mr L comes to the party, only \(n-k\) persons are present so he shakes hands with them

OpenStudy (rational):

adding \(\binom{n}{2}\) and \(n-k\) gives us the total handshakes which is \(100\)

OpenStudy (praxer):

I apply with this but I am confused though. please check it

OpenStudy (swamijatin2):

Let n be the number of people at the party. Each person shakes hands with every other person, so each person shakes hands with (n - 1) people, a possible total of n(n - 1) handshakes. But when person A shakes hands with person B, B also shakes hands with A, so each handshake would be counted twice. → number_of_handshakes = n(n - 1)/2 total number of handshakes is 100 → n(n - 1)/2 = 100 → n(n - 1) = 200 → n^2 - n - 200 = 0 As 200 is negative, one factor is positive and one is negative, so we need the factor pair of 200 which has a difference of 1, namely: 20 x 10 → (n - 20)(n + 10) = 0 → n = 20 or -10 There cannot be a negative number of people → there are 20 people present.

OpenStudy (rational):

@praxer your last line agrees with mine, but ur solution looks messy

OpenStudy (praxer):

yes it does. :(

OpenStudy (rational):

actually no, you/me have mistake somewhere. I am getting this : \[\large \binom{n}{2} + n-k = 100 \] \[\large \dfrac{n(n-1)}{2} + n-k = 100 \] \[\large n(n+1)= 2(k+100) \]

OpenStudy (swamijatin2):

so u get

OpenStudy (anonymous):

can you help me swamijatin

OpenStudy (anonymous):

in math on my question

OpenStudy (praxer):

@rational can you please explain your first proposition. I mean how you gave that. please explain :)

OpenStudy (anonymous):

me

OpenStudy (swamijatin2):

what is your queston @urbanmorgans

OpenStudy (anonymous):

in math first question

OpenStudy (anonymous):

can you help me

OpenStudy (swamijatin2):

i am not getting to question

OpenStudy (rational):

number of handshakes with \(n\) people is given by \(\large \binom{n}{2}\) you familiar with this right ?

OpenStudy (praxer):

yes :)

OpenStudy (anonymous):

why not it's called malinda

OpenStudy (swamijatin2):

Let n be the number of people at the party. Each person shakes hands with every other person, so each person shakes hands with (n - 1) people, a possible total of n(n - 1) handshakes. But when person A shakes hands with person B, B also shakes hands with A, so each handshake would be counted twice. → number_of_handshakes = n(n - 1)/2 total number of handshakes is 100 → n(n - 1)/2 = 100 → n(n - 1) = 200 → n^2 - n - 200 = 0 As 200 is negative, one factor is positive and one is negative, so we need the factor pair of 200 which has a difference of 1, namely: 20 x 10 → (n - 20)(n + 10) = 0 → n = 20 or -10 There cannot be a negative number of people → there are 20 people present.

OpenStudy (anonymous):

sorry melinda has saved

OpenStudy (rational):

\[\large n(n+1)= 2(k+100) \] By trial and error i am getting \(n=14\) and \(k=5\) so there are \(14\) people at the party excluding Mr L and \(5\) ppl have left before the arrival of Mr L

OpenStudy (praxer):

thank you :)

OpenStudy (rational):

yw